OpenStudy (anonymous):

2+(5)/(3k-6)=(-2)/(3k-6)^2

OpenStudy (anonymous):

\[k \approx2.299, -1.823\]

OpenStudy (anonymous):

They have to be fractions :[

OpenStudy (anonymous):

k≈2.299,−1.823

OpenStudy (anonymous):

Since we know that we are dealing with fractions, there are a couple ways to solve this: 1) Cross multiply. 2) Find a common denominator. 3) Multiply both sides with the variable with the biggest degree in the denominator to get rid of the fraction I find it easiest to eliminate the fraction since we are dealing with the same (3k-6) on one side but (3k-6)^2 on the other, meaning (3k-6) (3k-6). So let's go through this question using that method. Whichever one works best for you might be different. 2+(5)/(3k-6)=(-2)/(3k-6)^2 In this case, (3k-6)^2 would be the factor we multiply both sides with. (3k-6)^2 [2+(5)/(3k-6)] = (3k-6)^2 [(-2)/(3k-6)^2] Cancel out some of the fractions to simplify things a bit. (3k-6) (2+(5)) = (-2) (3k-6) (7) = (-2) (21k- 42) = (-2) 21k = (-2) + (42) 21k = (40) k = 40/21